TooTall, Inc. had dividends per share (DPS) of $2.00 on October 12, 2012. Seven years later, on October 12, 2019, DPS was $3.00. This growth trend is expected to continue into the indefinite future. If the required rate of return on TooTall stock is 18%, at what price should the stock sell today (October 13, 2019)?
Answer #1
1] Growth rate in dividends = (3/2)^(1/7)-1 = 5.96%
2] Stock price today = 3*1.0596/(0.18-0.0596) = $                 26.40